# The Hidden Power of Normalization: Exponential Capacity Control in Deep Neural Networks

### v1

@@ 4.3 Smoothening the loss landscape
Consider any member ${\bm{h}}_{no}\in{\mathcal{H}}_{no}$ defined in Definition 8. Denote $\left\|{\bm{h}}_{no},{\bm{W}}_{i}\right\|_{Lip}$ as the Lipschitz constant of ${\bm{h}}_{no}$ with respect to the weight ${\bm{W}}_{i}$, $\left\|{\bm{h}}_{no},{\bm{y}}_{i}\right\|_{Lip}$ as the Lipschitz constant of ${\bm{h}}_{no}$ w.r.t to the total input ${\bm{y}}_{i}={\bm{W}}_{i}{\bm{h}}_{i-1}$ at layer $i\leq K$, and $A_{i}=\sup_{{\bm{x}}}\|{\bm{h}}_{i}({\bm{x}})\|$ as the largest norm of the output of layer $i$. Then
@@
$\left\|{\bm{h}}_{no},{\bm{y}}_{i}\right\|_{Lip}$ $\leq$ $\begin{cases}\|{\mathsf{NO}}_{K}\|_{Lip}&\text{if }i=K\\ \|{\bm{W}}_{i+1}\|\cdots\|{\bm{W}}_{K}\|\prod_{k=i}^{K}\|{\mathsf{NO}}_{k}\|_{Lip}&\text{otherwise}\end{cases}$ (7)
This thoerem reveals that a normalization operator can have a direct control on the behaviors of many parts of a DNN. When the input variances at a layer is large, such a control will be stronger and a normalizer can significantly smoothen those parts of a DNN. Such a control should be crucial for the training process, since the training loss should be smooth in a controllable way. On the other side, an unnormalized network can be very bad as revealed below, whose proof appears in Appendix C.
@@
$\left\|f({\bm{h}}_{no},{\bm{x}}),{\bm{W}}_{i}\right\|_{Lip}\leq O(\|{\bm{W}}_{i+1}\|\cdots\|{\bm{W}}_{K}\|)\prod_{k=i}^{K}\|{\mathsf{NO}}_{k}\|_{Lip}$ (10)
For all $a_{j}\in[0,s_{j}]$ with $j>i$, there exists a member $\bm{h}$$^{*}\in{\mathcal{H}}$ satisfying:
$\left\|f({\bm{h}}^{*},{\bm{x}}),{\bm{W}}_{i}\right\|_{Lip}=a_{i+1}\cdots a_{K}\left\|\left({\bm{h}}_{i-1}({\bm{x}})\right)^{(n_{i+1},...,n_{K})}\right\|.$ (11)
@@
$\frac{\partial L(h^{*},{\bm{D}})}{\partial{\bm{W}}_{it}}=\frac{c{\color[rgb]{1,0,0}a_{K}\cdots a_{i+1}}}{m}\sum\limits_{{\bm{x}}\in{\bm{D}}}\frac{\partial(f\odot g)(u({\bm{x}}))}{\partial u}{\mathcal{I}}[{\bm{W}}_{it}{\bm{h}}_{i-1}({\bm{x}})\geq 0]({\bm{h}}_{i-1}({\bm{x}}))^{\top}$ (12)
for any $t\leq\min\{n_{i+1},...,n_{K}\}$ and $\frac{\partial L}{\partial{\bm{W}}_{it}}=0$ otherwise, where ${\bm{W}}_{it}$ is the $t$-th row of ${\bm{W}}_{i}$, $u({\bm{x}})={\bm{W}}{\bm{h}}_{K}({\bm{x}})$, and ${\mathcal{I}}(\cdot)$ is the indicator which is 1 if its input is true and 0 otherwise. $\odot$ denotes the function composition.
@@ 5.1 Benefits on training
Iteration complexity: Consider the problem of minimizing a function $\ell({\bm{w}})$. For convex problems, any first-order method requires at least $\Omega(\|\ell,{\bm{w}}\|_{Lip}^{2}/\alpha^{2})$ iterations to find an $\alpha$-approximate solution, while $\Omega(\|\ell,{\bm{w}}\|_{Lip}^{2}/\alpha)$ iterations are required for strongly convex problems (Bubeck et al., 2015). On the other hand, for nonsmooth nonconvex problems, Zhang et al. (2020) show that generalized gradient based methods need at least $\Omega(\|\ell,{\bm{w}}\|_{Lip}/\alpha)$ iterations to find an $\alpha$-stationary point. Combining those results with Corollary 14, the following result readily follows.
@@
For a DNN, the training loss is often nonconvex, Lipschitz, but sometimes nonsmooth, e.g. when using ReLU activations. Therefore this corollary points out that a learning algorithm can require a significant number of iterations to train an unnormalized model. The product $s_{i+1}\cdots s_{K}$ suggests that the weights at a deeper layer may require fewer interations to converge. In the worst case, the weight of the first layer can require $\Omega(s_{2}\cdots s_{K}/\alpha)$ iterations. Figure 1 suggests that $s_{2}\cdots s_{K}$ can be extremely huge in practice. As a result, training an unnormalized DNN can be exponentially expensive in the worse case.
@@
Consider the loss $L_{{\mathcal{H}}_{no}}$ and $L_{{\mathcal{H}}}$ being the empirical loss, defined for members of the family ${\mathcal{H}}_{no}$ and $\mathcal{H}$}, respectively, where $\mathsf{BN}$} is used. Let $\sigma$ be defined as in Corollary 10. Then for any $\alpha>0$, an $\alpha$-stationary point for the weights at layer $i<K$ can be found with
@@
When the input variances ($\sigma$) are large, this corollary suggests that a normalized network can be trained much more efficiently than its unnormalized version. There can be an exponential reduction ($\sigma^{-K+i-1}$) in the number of required interations, especially for the weights at some first layers of a DNN. This is a significant advantage of a normalizer.
@@ 5.2 Generalization ability of normalized DNNs
We next discuss the consequences about generalization ability of normalized models. As indicated before, a normalizer can make the Lipschitz constant of a model much smaller. This is an important capacity control. This section shows how Lipschitz control can translate into generalization improvement. Such a control can help a model generalize better.
@@ 5.2.1 Local Lipschitz continuity implies generalization
Consider the setting and notations in Theorem 20. Let ${\mathcal{X}}=\bigcup_{i=0}^{N}{\mathcal{X}}_{i}$ where ${\mathcal{X}}_{0}$ is the set of all instances so that the loss $f$ is not Lipschitz contiuous at any ${\bm{x}}\in{\mathcal{X}}_{0}$. If $P({\mathcal{X}}_{0})=0$ and $f$ is $L_{i}$-Lipschitz continuous w.r.t. ${\bm{x}}\in{\mathcal{X}}_{i}$ for any $i\in{\bm{T}},i>0$, then with probability at least $1-\delta$: $F(P,{\bm{h}})\leq F({\bm{D}},{\bm{h}})+\sum_{i\in{\bm{T}}}\frac{m_{i}}{m}\lambda_{i}L_{i}+g({\bm{D}},\delta).$
@@
$F(P,{\bm{h}})\leq F({\bm{D}},{\bm{h}})+\sum_{i\in{\bm{T}}}\frac{m_{i}}{m}\lambda_{i}\max\{1,L_{i}\}+\max_{i}\lambda_{i}L\sqrt{\frac{\ln(4/\delta)}{m}}+\sqrt{\frac{N}{m}}+const.,$ (HKK)
where $L=\max\{1,\|f,{\bm{x}}\|_{Lip}\}$, requiring the loss to be globally Lipschitz contiuous. As pointed out before, such an assumption makes this bound already more restrictive than ours. Furthermore, the term $\sqrt{N/m}$ causes their bound to surfer from the curse of dimensionality, since $N=\lambda^{-O(n)}$ in the worst case, where $n$ is the input dimensionality, meanwhile our bound does not. Those reasons suggest that the bound (HKK) is loose and significantly inferior to ours. Furthermore, their bound cannot be applied to the cases of discontinuity, while our result applies well to those cases as shown by Corollary 21.
@@ 5.2.2 Consequences for normalized DNNs
Consider a neural network $\bm{h}$} defined in Definition 4, and its normalized version ${\bm{h}}_{no}$. Given the assumption in Theorem 20 and notations in Corollary 10 and Corollary 11, assume further that loss $f({\bm{h}},\cdot)$ is $L_{f}$-Lipschitz continuous w.r.t $\bm{h}$}. Denoting $\omega({\bm{h}})=L_{f}P_{w}({\bm{h}})\sum_{i\in{\bm{T}}}\frac{m_{i}}{m}\lambda_{i}$, we have the followings with probability at least $1-\delta$:
@@
$F(P,{\bm{h}}_{no})-F({\bm{D}},{\bm{h}}_{no})$ $\leq g({\bm{D}},\delta)+\omega({\bm{h}}_{no})\sigma^{K}$ $(\text{for }{\mathsf{BN}})$ (15)
$F(P,{\bm{h}}_{no})-F({\bm{D}},{\bm{h}}_{no})$ $\leq g({\bm{D}},\delta)+\omega({\bm{h}}_{no})\prod_{k=1}^{K}(1-n_{k}^{-1})/{\sigma_{k}}$ $(\text{for }{\mathsf{LN}})$ (16)
This theorem suggests that $\mathsf{BN}$} and $\mathsf{LN}$} can significantly reduce the generalization error of a DNN. Without normalization, quantity $\omega({\bm{h}})$ can be of order $P_{w}({\bm{h}})$ which can be extremely large in practice. The bound in (15) can be tighten by replacing $\sigma^{-K}$ with $\prod_{k=1}^{K}\|1/{\bm{\sigma}_{k}}\|$ as discussed in Section 4. In practice, the product $\prod_{k=1}^{K}1/{\sigma_{k}}$ can be exponentially small, as pointed out in Subsection 4.2 and Figure 2, suggesting that normalization operators can significantly improve generalization ability of a normalized DNN.
@@
$\mathbb{E}_{{\bm{u}}}\left[F(P,{\bm{h}}({\bm{x}},{\bm{W}}+{\bm{u}}))\right]\leq F({\bm{D}},{\bm{h}})+\frac{\tau\mu^{2}}{2}+C\sqrt{\frac{|{\bm{W}}|\mu^{-2}+2\ln(m/\delta)}{m-1}}+const.,$ (LLA)
@@
In other work, Cisneros-Velarde et al. (2025) introduced a new bound that successfully demonstrates the generalization benefit of Weight Normalization (WN). Their analysis, however, is confined to a specific architecture where each layer’s normalized weight matrix has a Frobeneous norm of 1. While their resulting $O(\rho_{1}^{2}/\sqrt{m})$ bound (where $\rho_{1}$ is the radius of the last layer’s weight domain) effectively highlights WN’s advantage, it is not generalizable. It cannot be applied to other normalization methods. Our bound, in sharp contrast, is broadly applicable to many influential normalizers, as shown in Corollary 22.
Connection to great success: Normalized DNNs with great success often belong to overparameterization regime where the number of parameters in a model may far exceed the training size $m$. Those overparameterized models often have a small (or even zero) training error, do not overfit, but generalize well on unseen data (Zhang et al., 2021). An explanation can be derived from Corollary 22. Indeed, consider a normalized network ${\bm{h}}_{no}$ with a small training error, i.e., $F({\bm{D}},{\bm{h}}_{no})\approx 0$. Corollary 22 implies that its expected loss can be $F(P,{\bm{h}}_{no})=O(m^{-0.5})$. This loss suggests that a normalized network is significantly more data-efficient than unnormalized networks, since DNNs without any special assumption may have a generalization error of $O(m^{-1/(n+3)}\log m)$ (Bach, 2017), where $n$ is the input dimensionality. Corollary 22 suggests that a normalized network requires significantly less data than unnormalized networks in order to generalize well. Furthermore, while an unnormalized DNN surfers from the curse of dimensionality, its normalized version can avoid this curse owing to capacity reduction ($\prod_{k=1}^{K}\|{1}/{\bm{\sigma}_{k}}\|$) which is often significant in practice.
@@ A.1 Properties of Batch norm
By definition, we have $\partial{\mathsf{BN}}(x_{k},\epsilon)/\partial x_{k}=1/{\sqrt{\sigma_{k}^{2}+\epsilon}}$, while $\partial{\mathsf{BN}}(x_{k},\epsilon)/\partial x_{j}=0$ for any $j\neq k$. So $\left\|\frac{\partial{\mathsf{BN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}\right\|=\max_{1\leq k\leq n}|\partial{\mathsf{BN}}(x_{k},\epsilon)/\partial x_{k}|=\|1/{\bm{\sigma}}\|$. Therefore
@@ A.2 Properties of Layer norm
Given $\epsilon>0$, and ${\bm{x}}\in\mathbb{R}^{n}$, we have $\frac{\partial{\mathsf{LN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}=\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}({\bm{I}}-\frac{{\bm{y}}^{T}{\bm{y}}}{\|{\bm{y}}\|_{2}^{2}+n\epsilon})({\bm{I}}-\frac{1}{n}\bm{1})$, where ${\bm{y}}={\bm{x}}({\bm{I}}-\frac{1}{n}\bm{1})$, $\bm{1}$ is the $n\times n$ matrix of $1$’s, $\bm{I}$} is the $n\times n$ identity matrix.
Proof We can write the $i$-th element of LN as
${\mathsf{LN}}({\bm{x}},\epsilon)_{i}=\frac{y_{i}}{\sqrt{\frac{1}{n}\sum_{j=1}^{n}y_{j}^{2}+\epsilon}}$
Observe that
$\frac{\partial{\mathsf{LN}}({\bm{x}},\epsilon)_{i}}{\partial y_{t}}$ $=$ $\frac{\partial}{\partial y_{t}}\left(\frac{y_{i}}{\sqrt{\frac{1}{n}\sum_{j=1}^{n}y_{j}^{2}+\epsilon}}\right)$ (17)
$=$ $\frac{\delta_{it}{\sqrt{\frac{1}{n}\sum_{j=1}^{n}y_{j}^{2}+\epsilon}}-y_{i}\frac{y_{t}/n}{{\sqrt{\frac{1}{n}\sum_{j=1}^{n}y_{j}^{2}+\epsilon}}}}{\frac{1}{n}\sum_{j=1}^{n}y_{j}^{2}+\epsilon}$ (18)
$=$ $\frac{1}{\sqrt{\frac{1}{n}\sum_{j=1}^{n}y_{j}^{2}+\epsilon}}(\delta_{it}-\frac{y_{i}y_{t}}{\sum_{j=1}^{n}y_{j}^{2}+n\epsilon})$ (19)
$=$ $\frac{1}{\sqrt{\frac{1}{n}\|{\bm{y}}\|_{2}^{2}+\epsilon}}(\delta_{it}-\frac{y_{i}y_{t}}{\|{\bm{y}}\|_{2}^{2}+n\epsilon})$ (20)
$=$ $\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}(\delta_{it}-\frac{y_{i}y_{t}}{\|{\bm{y}}\|_{2}^{2}+n\epsilon})$ (21)
where $\delta_{it}=1$ for $i=t$ and $\delta_{it}=0$ for $i\neq t$. In matrix form, we have
$\frac{\partial{\mathsf{LN}}({\bm{x}},\epsilon)}{\partial{\bm{y}}}$ $=$ $\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}({\bm{I}}-\frac{{\bm{y}}^{T}{\bm{y}}}{\|{\bm{y}}\|_{2}^{2}+n\epsilon})$ (22)
Therefore
$\frac{\partial{\mathsf{LN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}$ $=$ $\frac{\partial{\mathsf{LN}}({\bm{x}},\epsilon)}{\partial{\bm{y}}}\frac{\partial{\bm{y}}}{\partial{\bm{x}}}$ (23)
$=$ $\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}({\bm{I}}-\frac{{\bm{y}}^{T}{\bm{y}}}{\|{\bm{y}}\|_{2}^{2}+n\epsilon})({\bm{I}}-\frac{1}{n}\bm{1})$ (24)
Proof [Proof of Lemma 2]
From Lemma 23 we observe that
@@
$\left\|\frac{\partial{\mathsf{LN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}\right\|$ $=$ $\left\|\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}({\bm{I}}-\frac{{\bm{y}}^{T}{\bm{y}}}{\|{\bm{y}}\|_{2}^{2}+n\epsilon})({\bm{I}}-\frac{1}{n}\bm{1})\right\|$ (25)
@@
$\leq$ $\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}\left\|({\bm{I}}-\frac{{\bm{y}}^{T}{\bm{y}}}{\|{\bm{y}}\|_{2}^{2}+n\epsilon})\right\|\left\|({\bm{I}}-\frac{1}{n}\bm{1})\right\|$ (26)

### v2

@@ 4.3 Smoothening the loss landscape
Consider any member ${\bm{h}}_{no}\in{\mathcal{H}}_{no}$ defined in Definition 8. Denote $\left\|{\bm{h}}_{no},{\bm{W}}_{i}\right\|_{Lip}$ as the Lipschitz constant of ${\bm{h}}_{no}$ with respect to the weight ${\bm{W}}_{i}$, $\left\|{\bm{h}}_{no},{\bm{y}}_{i}\right\|_{Lip}$ as the Lipschitz constant of ${\bm{h}}_{no}$ w.r.t to the total input ${\bm{y}}_{i}={\bm{W}}_{i}{\bm{h}}_{i-1}$ at layer $i\leq K$, $b_{i}=\|{\bm{W}}_{i}\|\cdots\|{\bm{W}}_{K}\|$ and $A_{i}=\sup_{{\bm{x}}}\|{\bm{h}}_{i}({\bm{x}})\|$ as the largest norm of the output of layer $i$. Then
@@
$\left\|{\bm{h}}_{no},{\bm{y}}_{i}\right\|_{Lip}\leq\begin{cases}\|{\mathsf{NO}}_{K}\|_{Lip}&\text{if }i=K\\ b_{i+1}\prod_{k=i}^{K}\|{\mathsf{NO}}_{k}\|_{Lip}&\text{otherwise}\end{cases}$ (7)
This theorem reveals that a normalization operator can have a direct control on the behaviors of many parts of a DNN. When the input variances at a layer is large, such a control will be stronger and a normalizer can significantly smoothen those parts of a DNN. Such a control should be crucial for the training process, since the training loss should be smooth in a controllable way. On the other side, an unnormalized network can be very bad as revealed below, whose proof appears in Appendix C.
@@
$\left\|f({\bm{h}}_{no},{\bm{x}}),{\bm{W}}_{i}\right\|_{Lip}\leq O(b_{i+1})\prod_{k=i}^{K}\|{\mathsf{NO}}_{k}\|_{Lip}$ (10)
where $b_{i}=\|{\bm{W}}_{i}\|\cdots\|{\bm{W}}_{K}\|$. For all $a_{j}\in[0,s_{j}]$ with $j>i$, there exists a member $\bm{h}$$^{*}\in{\mathcal{H}}$ satisfying:
$\left\|f({\bm{h}}^{*},{\bm{x}}),{\bm{W}}_{i}\right\|_{Lip}=a_{i+1}\cdots a_{K}\left\|\left({\bm{h}}_{i-1}({\bm{x}})\right)^{(n_{i+1},...,n_{K})}\right\|$ (11)
@@
$\frac{\partial L(h^{*},{\bm{D}})}{\partial{\bm{W}}_{it}}=\frac{c{\color[rgb]{1,0,0}a_{K}\cdots a_{i+1}}}{m}\sum\limits_{{\bm{x}}\in{\bm{D}}}{\mathcal{I}}_{it}\frac{\partial(f\circ g)}{\partial u}\mid_{u=u({\bm{x}})}({\bm{h}}_{i-1}({\bm{x}}))^{\top}$ (12)
for any $t\leq\min\{n_{i+1},...,n_{K}\}$ and $\frac{\partial L}{\partial{\bm{W}}_{it}}=0$ otherwise, where ${\bm{W}}_{it}$ is the $t$-th row of ${\bm{W}}_{i}$, $u({\bm{x}})={\bm{W}}{\bm{h}}_{K}({\bm{x}})$, $\circ$ denotes the function composition, and ${\mathcal{I}}_{it}=\begin{cases}1&\text{ if }{\bm{W}}_{it}{\bm{h}}_{i-1}({\bm{x}})\geq 0,\\ 0&\text{otherwise}.\end{cases}$
@@ 5.1 Benefits on training
Iteration complexity: Consider the problem of minimizing a function $\ell({\bm{w}})$. For convex problems, any first-order method requires at least $\Omega(\|\ell,{\bm{w}}\|_{Lip}^{2}/\alpha^{2})$ iterations to find an $\alpha$-approximate solution, while $\Omega(\|\ell,{\bm{w}}\|_{Lip}^{2}/\alpha)$ iterations are required for strongly convex problems (Bubeck and others, 2015). On the other hand, for nonsmooth nonconvex problems, Zhang et al. (2020) show that generalized gradient based methods need at least $\Omega(\|\ell,{\bm{w}}\|_{Lip}/\alpha)$ iterations to find an $\alpha$-stationary point. Combining those results with Corollary 14, the following result readily follows.
@@
For a DNN, the training loss is often nonconvex, Lipschitz, but sometimes nonsmooth, e.g. when using ReLU activations. Therefore this corollary points out that a learning algorithm can require a significant number of iterations to train an unnormalized model. The product $s_{i+1}\cdots s_{K}$ suggests that the weights at a deeper layer may require fewer iterations to converge. In the worst case, the weight of the first layer can require $\Omega(s_{2}\cdots s_{K}/\alpha)$ iterations. Figure 1 suggests that $s_{2}\cdots s_{K}$ can be extremely huge in practice. As a result, training an unnormalized DNN can be exponentially expensive in the worse case.
@@
Consider $L_{{\mathcal{H}}_{no}}$ and $L_{{\mathcal{H}}}$ being the empirical losses, defined for members of the family ${\mathcal{H}}_{no}$ and $\mathcal{H}$}, respectively, where $\mathsf{BN}$} is used. Let $\sigma$ be defined as in Corollary 10. Then for any $\alpha>0$, an $\alpha$-stationary point for the weights at layer $i<K$ can be found with
@@
When the input variances ($\sigma$) are large, this corollary suggests that a normalized network can be trained much more efficiently than its unnormalized version. There can be an exponential reduction ($\sigma^{-K+i-1}$) in the number of required iterations, especially for the weights at some first layers of a DNN. This is a significant advantage of a normalizer.
@@ 5.2 Generalization ability of normalized DNNs
We next discuss the consequences about generalization ability of normalized models. As indicated before, a normalizer can make the Lipschitz constant of a model much smaller. This is an important capacity control. This section shows why Lipschitz control can translate into generalization improvement.
@@ 5.2.1 Local Lipschitz continuity implies generalization
Consider the setting and notations in Theorem 20. Let ${\mathcal{X}}=\bigcup_{i=0}^{N}{\mathcal{X}}_{i}$ where ${\mathcal{X}}_{0}$ is the set of all instances so that the loss $f$ is not Lipschitz continuous at any ${\bm{x}}\in{\mathcal{X}}_{0}$. If $P({\mathcal{X}}_{0})=0$ and $f$ is $L_{i}$-Lipschitz continuous w.r.t. ${\bm{x}}\in{\mathcal{X}}_{i}$ for any $i\in{\bm{T}},i>0$, then with probability at least $1-\delta$: $F(P,{\bm{h}})\leq F({\bm{D}},{\bm{h}})+\sum_{i\in{\bm{T}}}\frac{m_{i}}{m}\lambda_{i}L_{i}+g({\bm{D}},\delta).$
@@
$F(P,{\bm{h}})\leq F({\bm{D}},{\bm{h}})+\gamma(L)+\sqrt{{N}/{m}}+const.,$ (HKK)
where $\gamma(L)=\sum_{i\in{\bm{T}}}\frac{m_{i}}{m}\lambda_{i}\max\{1,L_{i}\}+\max_{i}\lambda_{i}L\sqrt{\frac{\ln(4/\delta)}{m}}$ and $L=\max\{1,\|f,{\bm{x}}\|_{Lip}\}$, requiring the loss to be globally Lipschitz continuous. As pointed out before, such an assumption makes this bound already more restrictive than ours. Furthermore, the term $\sqrt{N/m}$ causes their bound to suffer from the curse of dimensionality, since $N=\lambda^{-O(n)}$ in the worst case, where $n$ is the input dimensionality, meanwhile our bound does not. Those reasons suggest that the bound (HKK) is loose and significantly inferior to ours. Furthermore, their bound cannot be applied to the cases of discontinuity, while our result applies well to those cases as shown by Corollary 21.
@@ 5.2.2 Consequences for normalized DNNs
Consider a neural network $\bm{h}$} defined in Definition 4, and its normalized version ${\bm{h}}_{no}$. Given the assumption in Theorem 20 and notations in Corollary 10 and Corollary 11, assume further that loss $f({\bm{h}},\cdot)$ is $L_{f}$-Lipschitz continuous w.r.t $\bm{h}$}. The followings hold with probability at least $1-\delta$:
@@
$F(P,{\bm{h}}_{no})-F({\bm{D}},{\bm{h}}_{no})$ $\leq$ $g({\bm{D}},\delta)+\omega({\bm{h}}_{no})\sigma^{-K}\;\;\;\;\;\;\;\;\;\;\;\;({\mathsf{BN}})$
$F(P,{\bm{h}}_{no})-F({\bm{D}},{\bm{h}}_{no})$ $\leq$ $g({\bm{D}},\delta)+\omega({\bm{h}}_{no})\prod_{k=1}^{K}1/{\sigma_{k}}\;\;\;\;\;({\mathsf{LN}})$
where $\omega({\bm{h}})=L_{f}P_{w}({\bm{h}})\sum_{i\in{\bm{T}}}\frac{m_{i}}{m}\lambda_{i}$.
This theorem suggests that $\mathsf{BN}$} and $\mathsf{LN}$} can significantly reduce the generalization error of a DNN. Without normalization, quantity $\omega({\bm{h}})$ can be of order $P_{w}({\bm{h}})$ which can be extremely large in practice. In practice, the product $\prod_{k=1}^{K}1/{\sigma_{k}}$ can be exponentially small, as pointed out in Subsection 4.2 and Figure 2, suggesting that normalization operators can significantly improve generalization ability of a normalized DNN.
@@
$\mathbb{E}_{{\bm{u}}}\left[F(P,{\bm{h}}({\bm{x}},{\bm{W}}+{\bm{u}}))\right]\leq F({\bm{D}},{\bm{h}})+\frac{\tau\mu^{2}}{2}+\frac{O(\ln m)}{\mu\sqrt{m}}+const.,$ (LLA)
@@
In other work, Cisneros-Velarde et al. (2025) introduced a new bound that successfully demonstrates the generalization benefit of Weight Normalization (WN). Their analysis, however, is confined to a specific architecture where each layer’s normalized weight matrix has a Frobenious norm of 1. While their resulting $O(\rho_{1}^{2}/\sqrt{m})$ bound (where $\rho_{1}$ is the radius of the last layer’s weight domain) effectively highlights WN’s advantage, it is not generalizable. It cannot be applied to other normalization methods. Our bound, in sharp contrast, is broadly applicable to many influential normalizers, as shown in Corollary 22.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
We would like to thank Bum Jun Kim for pointing out some typos or inconsistencies.
@@ A.1 Properties of Batch norm
By definition, we have $\partial{\mathsf{BN}}(x_{k},\epsilon)/\partial x_{k}=1/{\sqrt{\sigma_{k}^{2}+\epsilon}}$, while $\partial{\mathsf{BN}}(x_{k},\epsilon)/\partial x_{j}=0$ for any $j\neq k$. So $\left\|\frac{\partial{\mathsf{BN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}\right\|=\|1/{\bm{\sigma}}\|$. Therefore
@@ A.2 Properties of Layer norm
Given $\epsilon>0$, and ${\bm{x}}\in\mathbb{R}^{n}$, we have
@@
$\frac{\partial{\mathsf{LN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}=\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}\left({\bm{I}}-\frac{{\bm{y}}{\bm{y}}^{T}}{\|{\bm{y}}\|^{2}+n\epsilon}\right)\left({\bm{I}}-\frac{1}{n}\bm{1}\right)$
where ${\bm{y}}=({\bm{I}}-\frac{1}{n}\bm{1}){\bm{x}}$, $\bm{1}$ is the $n\times n$ matrix of $1$’s, and $\bm{I}$} is the identity matrix.
Proof We can explicitly write $y_{i}=x_{i}-\mu$ for the $i$-th element in $\bm{y}$}. The $i$-th output of LN is $z_{i}=\frac{y_{i}}{\sqrt{\sigma_{n}^{2}+\epsilon}}$. Recall that $\sigma_{n}^{2}+\epsilon=\frac{1}{n}\sum_{j}y_{j}^{2}+\epsilon=\frac{1}{n}\|{\bm{y}}\|^{2}+\epsilon$. The partial derivative with respect to the centered vector $\bm{y}$} is:
@@
$\frac{\partial z_{i}}{\partial y_{t}}$ $=$ $\frac{\delta_{it}\sqrt{\frac{1}{n}\|{\bm{y}}\|^{2}+\epsilon}-y_{i}\left(\frac{1}{2\sqrt{\frac{1}{n}\|{\bm{y}}\|^{2}+\epsilon}}\cdot\frac{2y_{t}}{n}\right)}{\frac{1}{n}\|{\bm{y}}\|^{2}+\epsilon}$ (15)
$=$ $\frac{1}{\sqrt{\sigma_{n}^{2}+\epsilon}}\left(\delta_{it}-\frac{y_{i}y_{t}}{n(\sigma_{n}^{2}+\epsilon)}\right)$ (16)
$=$ $\frac{1}{\sqrt{\sigma_{n}^{2}+\epsilon}}\left(\delta_{it}-\frac{y_{i}y_{t}}{\|{\bm{y}}\|^{2}+n\epsilon}\right)$ (17)
where $\delta_{it}=1$ for $i=t$ and $\delta_{it}=0$ for $i\neq t$. In matrix form (Jacobian w.r.t $\bm{y}$}):